HETAV PATEL

Grade 11,

3 capacitor having capacitance 8.4,8.2and 4.2microfard are connected in series across a 36V potential difference. (A)what is total energy stored in all 3 capacitor? (B)the capacitors are connected from potential difference without allowing them to discharge.they are reconnected in parallel with each other,with positively charged plates connected together. what is voltage across each capacitor in parallel combination? please give the whole solution.


























 

3 capacitor having capacitance 8.4,8.2and 4.2microfard are connected in series across a 36V potential difference.


(A)what is total energy stored in all 3 capacitor?


(B)the capacitors are connected from potential difference without allowing them to discharge.they are reconnected in parallel with each other,with positively charged plates connected together. what is voltage across each capacitor in parallel combination?


please give the whole solution.




Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shefali

capacitors are connected in series

so equivalent capacitance    1/C =1/C1 + 1/C2 + 1/C3

                                             =1/8.4  + 1/8.2 + 1/4.2

                                   C = 2.087 μF

 so Total energy stored U = 1/2 C V2

                                    = 1352.37  μJ

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin